OpenStudy (anonymous):

The first three digits are determined by the ZIP Code of the mailing address shown on the application for the Social Security number and the rest have restrictions against the middle two (group #) being 00 and the last four (serial #) being 0000. In Texas, the SSN may begin with 449 – 467 or 627 - 645. On June 25, 2011, Social Security changed the process to SSN Randomization to better protect the identity of each individual. how many SSN’s may be assigned in Texas?

OpenStudy (anonymous):

1. According to Social Security Number Verification Service since 1973 social security numbers have been assigned using the following method. The first three digits are determined by the ZIP Code of the mailing address shown on the application for the Social Security number and the rest have restrictions against the middle two (group #) being 00 and the last four (serial #) being 0000. In Texas, the SSN may begin with 449 – 467 or 627 - 645. On June 25, 2011, Social Security changed the process to SSN Randomization to better protect the identity of each individual. Using this information, how many SSN’s may be assigned in Texas? According to the US Census Bureau the population of Texas in 2009 was 24,782,302; will we run out of numbers? ABC = 449-467 or 627-645 DE = 01-99 FGHJ = 0001-9999 38 * 99 * 9999 = 35,636, 436, so not going to run out of numbers